# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 5 Rate of convergence
In order to evaluate the performance of Flow Matching, we are going to derive a rate of convergence using neural networks for the set $\mathcal{M}.$ Since $\tilde{v}$ is chosen such that it minimizes (8) instead of (4), we cannot use the minimization property in Theorem˜3.1 directly. The next theorem provides a more detailed decomposition of the error in the form of a classical oracle inequality. In order to merge the results later, we choose $\mu_{t}$ such that the results of Section˜4 apply. For the variance function $\sigma_{t}$, the analysis in Section˜4 reveals that we are not restricted to the linear case. Thus, we can choose a variance function that is suited for the application of a Bernstein-type inequality. Since Bernstein-type inequalities rely on absolute value bounds, we are going to choose $\sigma_{t}$ such that the absolute value of (13) is as small as possible. Our choice follows from Lemma˜4.3 and the solution of the ODE
$\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{{\min}}),\quad\sigma_{0}=1.$
@@
Let $p^{*}$ be of the form (14) and grant ˜3.3 and ˜5.1. Assume that $\log(\sigma_{\min})\sim\log(n)$. Then we have for every $a\in(0,1],\tau\in\mathbb{R}_{>0},b>1$ and $n$ big enough with probability of $1-\frac{1}{n}$ that
@@
$\quad+\frac{a^{-1}\log(n)^{7}}{n}\log\big{(}2\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\big{)}+(2+a)\tau+a\log(n)^{6},$
where $\tilde{v}\in\mathcal{M}$ is chosen such that it is arbitrarily close to realizing the infimum in (16).
The proof of Theorem˜5.2 uses the Bernstein-type concentration inequality of Chen et al. (2023b), which is commonly used in the analysis of diffusion models, see for example Yakovlev and Puchkin (2025). Theorem˜5.2 shows that a careful choice of the variance shift can allow for to a logarithmic dependency of the oracle bounds on $\sigma_{\min}^{-1}$ instead of a linear dependency.
In practice, the parameterized function $\hat{v}$ is typically represented by a neural network. To analyze the theoretical performance in this setting, we connect Theorem˜5.2 with known approximation results for neural networks. Specifically, we consider the following fully connected feedforward networks with ReLU activation functions, which have been extensively studied (see, e.g., Yarotsky (2017), Gühring et al. (2020), Kohler and Langer (2021), Schmidt-Hieber (2020), Suzuki (2019)).
The Rectified Linear Unit (ReLU) activation is defined as $\phi\colon\mathbb{R}\rightarrow\mathbb{R}$, with
$\phi(x)\coloneqq\max(0,x).$
For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, the shifted ReLU activation $\phi_{v}\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is defined componentwise by
@@
Using networks of the form (17) clipped at (15) for the set $\mathcal{M}$, the next result combines the findings of Section˜4, (12) and Theorem˜5.2 with approximation theory of Gühring et al. (2020) that allows for simultaneous approximation of a function and its derivative. To facilitate the transfer of results, we refrain from inserting $\alpha=1$, which is the applicable smoothness in this setting as shown in Lemma˜7.11.
@@
Let $p^{*}$ be of the form (14) and grant ˜3.3 and ˜5.1. Then for $n$ large enough with probability of $1-\frac{1}{n}$ for fixed $\eta>0$
@@
where $\hat{\psi}$ is the solution of an ODE whose vector field is in $\mathcal{M}$ given by a ReLU neural network, with no more than $c\cdot\log\left(n\right)$ layers, $c\cdot n^{c(d,\alpha,\eta)}\cdot\log^{2}\left(n\right)$ nonzero weights, where $c$ and $c(d,\alpha,\eta)$ is a constant independent of $n$.
[…]
Compared to Kunkel and Trabs (2025b, Theorem 3.6), our result applies to networks with logarithmically growing depth and polynomially growing numbers of non-zero weights. On the other hand, the rate deteriorates slightly. However, the two results cannot be compared directly. Kunkel and Trabs (2025b, Theorem 3.6) assume that the support of $p^{*}$ is compact, whereas Theorem˜5.3 assumes full support on $\mathbb{R}^{d}$. The evaluation metric differs from that in Gao et al. (2024b), who use the Wasserstein $2$ metric. According to Villani (2008, Remark 6.5), their rate $n^{-\frac{1}{d+5}}$ is also valid for the Wasserstein $1$ metric. For $\alpha\neq 0$ and large $d$, the rate obtained in Theorem˜5.3 is faster. However, this is just a rough comparison, as the unknown distributions studied are different. Furthermore, they use a linear variance shift and an early stopping approach, which hinders direct comparison even more.
@@ 6 Limitations and outlook
Our analysis provides a theoretical explanation for the empirical success of Flow Matching and aligns more closely with real-world scenarios than previous results.
Section˜4 paves the way for further research into broader classes of distributions whose vector fields have a bounded Lipschitz constant. The class of functions of form (14) served as a toy example and is an interesting starting point for generalization. Conversely, the lower bound on the Lipschitz constant could be used to identify distributions that cannot be mimicked with a "good" rate of convergence. Another promising direction is the introduction and consequences of an artificial Lipschitz control of the vector field. This influences the equivalence of gradients between (4) and (7) as well as their empirical counterparts. Controlling this effect could extend convergence results to larger classes of unknown distributions.
Finally, using specifically tailored approximation results or more sophisticated neural networks could improve the convergence rate further.
@@ 7 Proofs
7.1 Proofs of Section˜2
Proof of Lemma˜2.1.
Setting $\tilde{v}_{t}=v_{t}$, we obtain
$0=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}+C\quad\Longleftrightarrow\quad C=-\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}.$
This concludes the proof. ∎
7.2 Proofs of Section˜4
Proof of Theorem˜4.2.
First we show that we can calculate the Jacobian of $v$ explicitly. The proof and all subsequent proofs of auxiliary results are deferred to Section˜7.4.
@@ 7.2 Proofs of Section˜4
By assumption $\operatorname{Cov}(\operatorname{Cov}(Y^{x,t})_{ji})$ is bounded for all $x$. For the upper bound we use that by the mean value theorem, there exists an $\xi\in\mathbb{R}^{d}$ such that for the $i$-th coordinate function $v_{t}^{j}$
@@
with
$\lim_{|h|\rightarrow 0}\frac{|r(h)|}{|h|}=0.$
Let the smallest Lipschitz constant of $v_{t}$ be $\Gamma_{t}$. Using (19) we can conclude
$|D_{x}v_{t}(a)h|=|v_{t}(a+h)-v_{t}(a)-r(h)|$ $\leq|v_{t}(a+h)-v_{t}(a)|+|r(h)|$
$\leq\Gamma_{t}|h|+|r(h)|.$
@@
Since $i,j$ were arbitrary and the entries of the Jacobian are of the form (18), we obtain the result. ∎
Proof of Lemma˜4.3.
  1. 1.
    We begin by setting
    $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=h_{t},$
    where $h_{t}$ is a continuous function on $[0,1].$ By separations of variables, all of the solutions of this ODE are of the form
    $\sigma_{t}=ce^{H_{t}},$
    where $H_{t}$ is an anti-derivative of $h_{t}$ and $c\in\mathbb{R}$. we use $\sigma_{0}=1$ as initial condition, which leads to
    $1=ce^{H_{0}}\quad\Longleftrightarrow\quad c=\frac{1}{e^{H_{0}}}.$
    To assure $\sigma_{1}=\sigma_{\min}$, we need to choose $H$ such that
    $\sigma_{\min}=\frac{e^{H_{1}}}{e^{H_{0}}}\quad\Longleftrightarrow\quad H_{1}-H_{0}=\log(\sigma_{\min}).$
    By the mean-value-theorem there is a $t^{*}\in[0,1]$ such that $H_{t^{*}}^{\prime}=h_{t^{*}}=\log(\sigma_{\min})$.
  2. 2.
    By change of variables we have that
    $\int_{0}^{1}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t=-\int_{0}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\;\mathrm{d}t=\int_{\sigma_{\min}}^{1}\frac{1}{u}\;\mathrm{d}u=\log(\sigma_{\min}^{-1}).$
Proof of Theorem˜4.5.
@@
Let $t^{*}$ be such that $\sigma_{t^{*}}=\frac{1}{\vartheta}$ for $\vartheta\in\mathbb{R}_{\geq 1}$. Grant ˜4.4 (III). Then
@@
For large $t$, we need to assume that $\operatorname{Cov}(Y^{x,t})_{ij}$ decays fast enough for $i\neq j$ and $t\rightarrow 1$ to bound the integral over all $B_{ij}$ for $i\neq j$. Under ˜4.4 (I), we know that for all $i\neq j$ and all $x\in\mathbb{R}^{d}$
$\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})$ $\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ij}|\;\mathrm{d}t\leq(t^{*})^{-2\gamma}\int_{t^{*}}^{1}\big{|}(\gamma t^{\gamma-1}\sigma_{t}-\sigma_{t}^{\prime}t^{\gamma})\big{|}\;\mathrm{d}t$
@@
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})|\;\mathrm{d}t$
we need to use ˜4.4 (II). Inserting the expression for the variance $\operatorname{Var}(Y^{x,t}_{i})$ we obtain
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})|\;\mathrm{d}t$ $=\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}\big{(}1+O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)})|\;\mathrm{d}t$
$\leq\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}|\;\mathrm{d}t$
$\quad+\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t.$
@@
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}|\;\mathrm{d}t$ $=\int_{t^{*}}^{1}|\gamma t^{-1}|\;\mathrm{d}t=-\gamma\log(t^{*}).$

### v2

@@ 4 Rate of convergence
The goal of this chapter is to derive a rate of convergence, e.g. an upper bound on $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})$ depending on the number of observations $n$, using the estimator $\hat{\psi}$ from (6). For the set $\mathcal{M}$ we are going to use a subclass of ReLU networks, which will be defined below. Leveraging the insights from Section˜3, we consider unknown distributions $\mathbb{P}^{*}$ that admit Lipschitz controlled vector fields and a noise schedule $\sigma_{t}$ that stabilizes $v$ over all $t\in[0,1]$. We recover a classical trade-off in $\sigma_{\min}$, balancing this tuning parameter leads to the rate in Theorem˜4.3.
We start from Theorem˜2.1: In case the unknown distribution is such that ˜3.4 holds and $p^{*}\in B^{\alpha}_{1,\infty}$, we can bound using (9)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\lesssim\sigma_{\min}^{1+\alpha}+\Big(\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big[|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big]\Big)^{1/2}.$ (14)
Since $\tilde{v}$ is chosen via (6) instead of (5), we cannot use the minimization property in this bound directly. Furthermore, the stability of $v_{t}$ depends on $\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|$, which deteriorates if $\sigma_{\min}$ is smaller. Thus, we expect a typical variance-bias trade-off in $\sigma_{\min}$ in (14). Note that the approach is different from the early-stopping approach natural to diffusions and classically adopted in Flow Matching analysis. Typically, instead of choosing $\sigma_{\min}>0$, this tuning parameter is set to 0 and an early stopping time is imposed. While the two approaches are equivalent regarding the variance, the early stopping imposes an additional bias. This is commonly addressed with a very small early stopping time, see e.g. Gao et al. (2024b, Lemma 4.3), Fukumizu et al. (2025, Lemma 11), which prevents the bound from capturing smoothness at this instance. The next theorem provides a more detailed decomposition of the former error in (14) providing a classical classical oracle inequality. In order to merge the results later, we choose $\mu_{t}$ such that the results of Section˜3 apply. For the variance function $\sigma_{t}$, the analysis in Section˜3 reveals that we are not restricted to the linear case. Thus, we can choose a variance function that is suited for the application of a Bernstein-type inequality. Since Bernstein-type inequalities rely on absolute value bounds, we are going to choose $\sigma_{t}$ such that the absolute value of (10) is as small as possible. Our choice follows from Lemma˜3.3 and the solution of the ODE $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{{\min}})$ with the initial condition $\sigma_{0}=1.$
@@
While the choice of $\sigma_{t}$ in this work is mainly for concentration convenience and to exploit the generality of Section˜3, Tsimpos et al. (2025) show optimality for similar schedules in dynamic transport. Given the variety of concentration bounds available for linear and diffusion-type schedules, see Gao et al. (2024b, Lemma 5.10), Azangulov et al. (2024, Theorem 34), Stéphanovitch et al. (2025, Proposition 12), the below analysis can be extended to other choices of $\sigma_{t}$.
@@
Let $p^{*}$ be of the form (13) and grant ˜2.3 and ˜4.1. Assume that $\log(\sigma_{\min}^{-1})\sim\log(n)$. Then we have for every $a\in(0,1],\tau\in\mathbb{R}_{>0},b>1$ and $n$ big enough with probability of $1-\frac{1}{n}$ that
@@
$\quad+\frac{a^{-1}\log(n)^{7}}{n}\log\big(2\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\big)+(2+a)\tau+a\log(n)^{6}.$
The proof of Theorem˜4.2 uses the Bernstein-type concentration inequality of Chen et al. (2023b), which is commonly used in the analysis of diffusion models, see for example Yakovlev and Puchkin (2025). Theorem˜4.2 shows that a careful choice of the variance shift can allow for to a logarithmic dependency of the oracle bounds on $\sigma_{\min}^{-1}$ instead of a linear dependency.
In practice, the parameterized function $\hat{v}$ is typically represented by a neural network. To analyze the theoretical performance in this setting, we connect Theorem˜4.2 with known approximation results for neural networks. Specifically, we consider the following fully connected feedforward networks with ReLU activation functions, which have been extensively studied, see, e.g., Yarotsky (2017), Gühring et al. (2020), Kohler and Langer (2021), Schmidt-Hieber (2020), Suzuki (2019).
The Rectified Linear Unit (ReLU) activation is defined as $\phi\colon\mathbb{R}\rightarrow\mathbb{R}$, with $\phi(x)\coloneqq\max(0,x).$ For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, the shifted ReLU activation $\phi_{v}\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is defined by
@@
Using networks of the form (17) clipped at (15) for the set $\mathcal{M}$, the next result combines the findings of Section˜3, (9) and Theorem˜4.2 with approximation theory of Gühring et al. (2020) that allows for simultaneous approximation of a function and its derivative. To facilitate the transfer of results, we refrain from inserting $\alpha=1$, which is the applicable smoothness in this setting as shown in Lemma˜A.3.
@@
Let $p^{*}$ be of the form (13) and grant ˜2.3 and ˜4.1. Then for $n$ large enough with probability of $1-\frac{1}{n}$ for fixed $\eta>0$
@@
where $\hat{\psi}$ is the solution of an ODE whose vector field is given by a ReLU neural network, with no more than $c\cdot\log\left(n\right)$ layers, $c\cdot n^{c(d,\alpha,\eta)}\cdot\log^{2}\left(n\right)$ nonzero weights, where $c$ and $c(d,\alpha,\eta)$ is a constant independent of $n$.
The proof of Theorem˜4.3 exploits the fact that the vector field $v$ from (4) is by construction in $C^{\infty}$. In order to bound the higher order derivatives of $v$, we leverage the fact that densities of the form (13) satisfy a $\log$-Sobolev inequality with a controlled constant. Interestingly, the dependence on the variance of $q$ from Theorem˜3.2 extends to a dependence on joint cumulants in higher order derivatives. The rate in Theorem˜4.3 benefits from smoothness in the unknown distribution. Since smoothing is an intrinsic property of Flow Matching, it is desirable that the rate also reflects this property. Even though the rate is not optimal, the gap to the optimal rate of $n^{-\frac{1+\alpha}{2\alpha+d}}$ is small for large $d$. It should be noted that this gap primarily results from the use of the general approximation result from Gühring et al. (2020), which bounds the supremum norm error and not the weighted $L_{2}$ error appearing in (16). This route reveals interesting properties of $v$, keeps the proof short and the paper focused on the distributional aspects of Flow Matching. A carefully fine tuned approximation result could close this gap.
Compared to Kunkel and Trabs (2025b, Theorem 3.6), our result applies to networks with logarithmically growing depth and polynomially growing numbers of non-zero weights. This aligns the setting of the theorem more closely with practical applications, which is necessary to explain the empirical success of Flow Matching. This comes at the cost of a much stricter assumption on the unknown distribution and, at least when refraining from an tailored network construction, a slightly suboptimal rate. Compared to Gao et al. (2024b), we gain in the rate since we are able to capture smoothness in $p^{*}$, while preserving a clean citation of the approximation result, but lose in the strictness of the evaluation metric ($\mathsf{W}_{1}$ instead of $\mathsf{W}_{2}$).
It should however be noted, that the three results cannot be compared directly. Kunkel and Trabs (2025b, Theorem 3.6) assume that the support of $p^{*}$ is compact, whereas Theorem˜4.3 assumes full support on $\mathbb{R}^{d}$. The distributions studied in Gao et al. (2024b) have full support, but the exact assumptions differ as well. Furthermore, they use a linear variance shift and an early stopping approach, which hinders direct comparison even more.
@@ 5 Limitations and outlook
Our analysis provides a theoretical explanation for the empirical success of Flow Matching and aligns more closely with real-world scenarios than previous results. Section˜3 paves the way for further research into broader classes of distributions whose vector fields have a bounded Lipschitz constant. The class of functions of form (12) and (13) served as a toy example and is an interesting starting point for generalization. Conversely, the lower bound on the Lipschitz constant could be used to identify distributions that cannot be mimicked with a "good" rate of convergence, like indicated in Example˜3.7. While ˜3.4 holds for a very broad range of variance functions, proofs based on concentration inequalities, such as Theorem˜4.3, depend on the specific choice of the variance function. Given optimal pushforward mappings, Tsimpos et al. (2025) have investigated optimal noise schedules. Interestingly, their optimal schedule connects to the choice in ˜4.1. Exploring optimal choices of variance functions from a statistical perspective would thus be very interesting. Another promising direction is the introduction and consequences of an artificial Lipschitz control of the vector field. This influences the equality in (5) as well as in the empirical counterpart. Controlling this effect could extend convergence results to larger classes of unknown distributions. A different, but interesting, question is whether there are Lipschitz controlled networks that approximate non-controlled vector fields and achieve good results. This requires a completely different approach in approximation theory. Another interesting aspect are bounds in different metrics. The bounds on the Jacobian in this work can serve as a starting point for analysis in the total variation distance, elaborating the ideas of Su et al. (2025) combined with Cai and Li (2025); Li et al. (2024).
@@ 6 Proofs
6.1 Proofs of Section˜3
Proof of Theorem˜3.2.
First we show that we can calculate the Jacobian of $v$ explicitly. The proof and all subsequent proofs of auxiliary results are deferred to Appendix˜A.
@@ 6.1 Proofs of Section˜3
By assumption $\operatorname{Cov}(Y^{x,t})_{ji}$ is bounded for all $x$. For the upper bound we use that by the mean value theorem, there exists an $\xi\in\mathbb{R}^{d}$ such that for the $i$-th coordinate function $v_{t}^{j}$
@@
with $\lim_{|h|\rightarrow 0}\frac{|r(h)|}{|h|}=0.$ Let the smallest Lipschitz constant of $v_{t}$ be $\Gamma_{t}$. Using (19) we can conclude
$|D_{x}v_{t}(a)h|=|v_{t}(a+h)-v_{t}(a)-r(h)|\leq|v_{t}(a+h)-v_{t}(a)|+|r(h)|\\ \leq\Gamma_{t}|h|+|r(h)|.$
@@
Since $i,j$ were arbitrary and the entries of the Jacobian are of the form (18), we obtain the bound on $\Gamma_{t}$.
For the existence of a $t^{*}$ such that $\frac{\sigma_{t^{*}}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min}^{-1})$, we begin by setting $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=h_{t},$ where $h_{t}$ is a continuous function on $[0,1].$ By separations of variables, all of the solutions of this ODE are of the form $\sigma_{t}=ce^{H_{t}},$where $H_{t}$ is an anti-derivative of $h_{t}$ and $c\in\mathbb{R}$. we use $\sigma_{0}=1$ as initial condition, which leads to $1=ce^{H_{0}}$ which is equivalent to $c=\frac{1}{e^{H_{0}}}.$ To assure $\sigma_{1}=\sigma_{\min}$, we need to choose $H$ such that $\sigma_{\min}=\frac{e^{H_{1}}}{e^{H_{0}}}$ which is equivalent to $H_{1}-H_{0}=\log(\sigma_{\min}).$ By the mean-value-theorem there is a $t^{*}\in[0,1]$ such that $H_{t^{*}}^{\prime}=h_{t^{*}}=\log(\sigma_{\min})$. ∎
Proof of Lemma˜3.3.
By change of variables we have that $\int_{0}^{1}\big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big|\;\mathrm{d}t=-\int_{0}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\;\mathrm{d}t=\int_{\sigma_{\min}}^{1}\frac{1}{u}\;\mathrm{d}u=\log(\sigma_{\min}^{-1}).$
Proof of Theorem˜3.5.
@@
Let $t^{*}$ be such that $\sigma_{t^{*}}=\frac{1}{\vartheta}$ for $\vartheta\in\mathbb{R}_{\geq 1}$. Grant ˜3.4 (III). Then
@@
For large $t$, we need to assume that $\operatorname{Cov}(Y^{x,t})_{ij}$ decays fast enough for $i\neq j$ and $t\rightarrow 1$ to bound the integral over all $B_{ij}$ for $i\neq j$. Under ˜3.4 (I), we know that for all $i\neq j$ and all $x\in\mathbb{R}^{d}$
$\int_{t^{*}}^{1}\Big|\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)$ $\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ij}\Big|\;\mathrm{d}t\leq(t^{*})^{-2\gamma}\int_{t^{*}}^{1}\big|(\gamma t^{\gamma-1}\sigma_{t}-\sigma_{t}^{\prime}t^{\gamma})\big|\;\mathrm{d}t$
@@
$\int_{t^{*}}^{1}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})\Big|\;\mathrm{d}t$
we need to use ˜3.4 (II). Inserting the expression for the variance $\operatorname{Var}(Y^{x,t}_{i})$ we obtain
$\int_{t^{*}}^{1}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})\Big|\;\mathrm{d}t=\int_{t^{*}}^{1}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}\Big(1+O\Big(\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{\frac{1}{\kappa}}\Big)\Big)\Big|\;\mathrm{d}t$
$\leq\int_{t^{*}}^{1}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}\Big|\;\mathrm{d}t+\int_{t^{*}}^{1}\Big|\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}O\Big(\Big(\frac{\sigma_{t}}{t^{\gamma}}\big)^{\frac{1}{\kappa}}\Big)\Big|\;\mathrm{d}t.$
@@
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+\Big(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\Big)\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big(\frac{\sigma_{t}}{t^{\gamma}}\Big)^{2}|\;\mathrm{d}t=\int_{t^{*}}^{1}|\gamma t^{-1}|\;\mathrm{d}t=-\gamma\log(t^{*}).$